OpenStudy (nuccioreggie):

How many solutions does the following equation have? |3x + 12| = 18 No solution One solution Two solutions Infinitely many solutions

OpenStudy (nuccioreggie):

@Michele_Laino

OpenStudy (michele_laino):

we have to distinguish two cases, namely such equation is equivalent to these systems: \[\Large \left\{ \begin{gathered} 3x + 12 = 18 \hfill \\ \hfill \\ 3x + 12 \geqslant 0 \hfill \\ \end{gathered} \right. \cup \left\{ \begin{gathered} - \left( {3x + 12} \right) = 18 \hfill \\ \hfill \\ 3x + 12 < 0 \hfill \\ \end{gathered} \right.\]

OpenStudy (nuccioreggie):

So its B

OpenStudy (michele_laino):

please wait, I'm doing some computation

OpenStudy (michele_laino):

no, I'm sorry, it is not option B

OpenStudy (nuccioreggie):

Oh wait i ment C

OpenStudy (michele_laino):

that's right!

OpenStudy (michele_laino):

from my computation, I got two distinct solutions